An AC circuit has a resistance of 500 ohms, a voltage of 200 V, an inductance of 0.4 mHz, a capacitance of 100pF and an angular frequency of 5.00 x 105 rad/s.

What is the impedance?

What is the current amplitude?

What is the voltage amplitude read by a voltmeter across the inductor, the resistor, and the capacitor individually?

What is the voltage amplitude read by a voltmeter across the inductor and capacitor together?
Answer #1

A] Reactance across the Inductor, XL = 2*pi*f*L = 2*3.14*5*105*0.4*10-3

XL = 1.2566 k-ohm

Reactance across the Capacitor, XC =1 / [2*pi*f*C] = 1 / [2*3.14*5*105*100*10-12] = 1 / [3.1416*10-4]

XC = 3.1831 k-ohm

Total Impedance, Z = sqrt [ R2 + (XC - XL)2] = sqrt [ (500)2 + (3.1831k - 1.2566k)2] = sqrt[3.96*106]

Total Impedance, Z = 1.990 k-ohm

B] Current amplitude, I = V/Z = 200 / 1.99*103 = 0.1 A